# Computing Exact Shapley Values in Polynomial Time for Product-Kernel Methods

### v1

@@ Abstract
Kernel methods are widely used in machine learning due to their flexibility and expressive power. However, their black-box nature poses significant challenges to interpretability, limiting their adoption in high-stakes applications. Shapley value-based feature attribution techniques, such as SHAP and kernel-specific variants like RKHS-SHAP, offer a promising path toward explainability. Yet, computing exact Shapley values remains computationally intractable in general, motivating the development of various approximation schemes. In this work, we introduce PKeX-Shapley, a novel algorithm that utilizes the multiplicative structure of product kernels to enable the exact computation of Shapley values in polynomial time. We show that product-kernel models admit a functional decomposition that allows for a recursive formulation of Shapley values. This decomposition not only yields computational efficiency but also enhances interpretability in kernel-based learning. We also demonstrate how our framework can be generalized to explain kernel-based statistical discrepancies such as the Maximum Mean Discrepancy (MMD) and the Hilbert-Schmidt Independence Criterion (HSIC), thus offering new tools for interpretable statistical inference.
1 Introduction
@@
Shapley values [1], a solution concept originating from cooperative game theory, offer a principled, axiomatic framework for feature attribution in machine learning (ML) [2, 3, 4]. Thanks to their rigorous axiomatic foundation, there has been a widespread adoption within explainable AI [5]. They allow a model’s output—such as predictions or losses—to be fairly distributed across input features based on their individual and joint contributions. Consequently, several methods have been proposed to estimate Shapley values, ranging from general model-agnostic methods like Kernel SHAP [2] to model-specific methods that leverage structural properties to improve statistical or computational efficiency. Well-known examples of the latter include Tree SHAP for tree-based models [6], GPSHAP for Gaussian processes [7], Deep SHAP for deep networks [2], and, most relevant to our work, RKHS-SHAP for kernel methods [8]. Kernel methods are particularly notable—not only for their use in prediction tasks but also in a broad range of statistical inference problems, including two-sample testing [9, 10, 11], goodness-of-fit testing [12, 13], (conditional) independence testing [14], causal inference and discovery [15, 16, 17, 18], among others. Hence, as kernel methods gain widespread adoption in high-stakes applications for their flexibility and expressive power, the need for interpretability has become increasingly vital.
@@
Despite their desirable properties, adopting Shapley values in machine learning generally faces two key challenges. The first is defining and estimating a suitable value function that quantifies the contribution of a coalition of features. Ideally, this function should reflect the model’s behavior when the complementary features are absent. A natural approach is to retrain the model on each subset and use the resulting prediction as the value [19]—an approach that is computationally infeasible due to the exponential number of retrainings required. A popular alternative simulates feature absence via marginal or conditional expectations of the model’s output, conditioned on the subset being fixed (see Sundararajan and Najmi [5] for other candidates). However, accurately estimating these expectations is difficult, often requiring density estimation or making simplifying assumptions such as feature independence, which often result in misleading and unfaithful explanations [20]. Chau et al. [8] addresses this under the context of kernel methods by utilizing the structure of reproducing kernel Hilbert spaces (RKHS) and adopting kernel distributional embeddings [21] to estimate value functions nonparametrically, thereby avoiding density estimation and independence assumptions.
@@
Given a value function, the second challenge is efficiently estimating the Shapley values. Exact computation requires evaluating the value function over all $2^{d}$ possible feature subsets for $d$ features, which is computationally infeasible. To address this, approximation techniques—such as Monte Carlo sampling and regression-based methods—are commonly used to estimate Shapley values from a smaller set of computed values. While these methods reduce computational costs, they introduce estimation errors that scale with dataset size and feature dimensionality [20]. Model-specific structures can sometimes mitigate this, as in Tree SHAP, which leverages tree decompositions for efficient computation. RKHS-SHAP, however, focuses on improving the statistical estimation of the value function, while Shapley values themselves are still approximated via regression-based methods—thus retaining the limitations mentioned above. Moreover, Chau et al. [8] focuses solely on learning algorithms and does not address the explanation of non-parametric statistical discrepancies, where kernel methods play a central role.
To enhance Shapley value-based attribution for kernel methods, we focus on a subclass that employs product kernels—referred to as product-kernel methods—and introduce PKeX-Shapley (Product-Kernel-based eXact Shapley attribution). Product kernels are widely used due to their computational efficiency and strong theoretical guarantees. For example, the product of universal kernels remains universal, meaning the resulting RKHS can approximate any continuous function, provided each component kernel is sufficiently expressive [22]. While our focus on product kernels narrows the scope relative to Chau et al. [8], which considers generic kernels, we show that this added structure brings significant benefits in addressing the two aforementioned challenges. First, we demonstrate that product-kernel methods naturally admit a functional decomposition—an additive sum over subsets of features. This motivates our choice of value function of set $\mathcal{S}$: we accumulate all components $\mathcal{S}^{\prime}$ from the functional decomposition for all $\mathcal{S}^{\prime}\subseteq\mathcal{S}$, following ideas similar to those in [23]. Second, leveraging this decomposition, we develop a recursive algorithm that computes exact Shapley values in polynomial time—eliminating the need for sampling or regression-based estimation, and significantly improving computational efficiency over existing methods. Another independent yet important contribution of our work is that we apply our computationally efficient framework to explain kernel-based statistical discrepancies such as the Maximum Mean Discrepancy (MMD) and the Hilbert space Independence Criterion (HSIC), leading to interpretable statistical inference. The Python implementation of this work is available at [24].
2 Background
@@
Notation. Let $\mathcal{D}$ denote the set of $d$ features, and $2^{\mathcal{D}}$ its power set. The training set ${(\boldsymbol{x}^{(i)},y^{(i)})}_{i=1}^{n}$ consists of $n$ samples, where $\boldsymbol{x}^{(i)}\in\mathbb{R}^{d}$ and $y^{(i)}\in\mathbb{R}$ (or a discrete label set for classification tasks). Let $\mathbf{X}\in\mathbb{R}^{n\times d}$ be the matrix of features, and $\mathbf{X}_{\mathcal{S}}\in\mathbb{R}^{n\times s}$ the submatrix restricted to features in subset $\mathcal{S}\subseteq\mathcal{D}$, and we write $\mathbf{X}_{j}:=\mathbf{X}_{\{j\}}$. We use capital letters for random variables, bold capital letters for matrices, calligraphic letters for sets, and bold lowercase letters for vectors. The restriction of a vector $\boldsymbol{x}$ to features in $\mathcal{S}$ is denoted by $\boldsymbol{x}_{\mathcal{S}}$. The elementwise product is denoted by $\odot$, and expectation by $\mathbb{E}$. A symmetric positive (semi-)definite kernel function over $\mathcal{D}$ is denoted by $k$, and its restriction to subset $\mathcal{S}$ by $k_{\mathcal{S}}$. The corresponding kernel matrices are denoted by $\mathbf{K}$ and $\mathbf{K}_{\mathcal{S}}$, respectively. All proofs are provided in Appendix C.
@@
Shapley value. The Shapley value [1], rooted in cooperative game theory, is a popular attribution method for computing feature importance in predictive models. It fairly distributes a model’s output across its input features by (weighted) averaging over their marginal contributions across all possible feature subsets. Specifically, given a value function (or characteristic function) $v:2^{\mathcal{D}}\to\mathbb{R}$, which quantifies the contribution of a feature subset, the Shapley value for a feature $j$ is defined as
@@
where $\mu(s)=\nicefrac{{s!(d-s-1)!}}{{d!}}$ and $v(\mathcal{S}\cup\{j\})-v(\mathcal{S})$ measures the marginal contribution of feature $j$ when added to subset $\mathcal{S}$. The Shapley value is the unique solution satisfying four axioms of efficiency, null player, symmetry, and linearity for any specified value function [1]. However, in the context of explainability, defining this value function $v$—which determines how subsets of features contribute to the model’s output—is itself a non-trivial challenge. Several definitions have been proposed [5], but two are widely adopted: the interventional value function $v_{\boldsymbol{x}}(\mathcal{S})=\mathbb{E}_{X_{\mathcal{D}\setminus\mathcal{S}}|X_{\mathcal{S}}=\boldsymbol{x}_{\mathcal{S}}}[f(\boldsymbol{x}_{\mathcal{S}},X_{\mathcal{D}\setminus\mathcal{S}})]$, which replaces missing features with samples from their marginal distributions [25, 26], and the observational value function $v_{\boldsymbol{x}}(\mathcal{S})=\mathbb{E}_{X_{\mathcal{D}\setminus\mathcal{S}}}[f(X)\,|\,X_{\mathcal{S}}=\boldsymbol{x}_{\mathcal{S}}]$, which imputes them using the conditional distribution given observed features [2]. While the observational approach preserves feature dependencies, it requires estimating complex conditional expectations, making it computationally demanding.
Many existing methods estimate value functions by sampling from marginal or conditional distributions, requiring repeated model evaluations that become costly for large models or high-dimensional data [5]. These estimates typically rely on a set of background samples (i.e., a subset of the training data), which can significantly influence the resulting explanations [27, Chapter 21]. Recent kernel-based approaches offer a more (statistically) efficient alternative by using kernel embeddings of marginal or conditional distributions [21] to derive closed-form expressions for nonparametric estimations of value functions in RKHS [8, 7]. While this formulation accelerates value function estimation, the Shapley value itself is still approximated via regression over a subset of feature coalitions, based on the Kernel SHAP [2] formulation.
In this work, we bypass the challenge of estimating value functions by adopting a functional decomposition approach—well-established in classical statistics, particularly in global sensitivity analysis [28] and analysis of variance [29, 30].
Functional decomposition. The functional decomposition of a function $f$ expresses it as a sum of components indexed by subsets of input variables: $f(\boldsymbol{x})=\sum_{\mathcal{S}\subseteq\mathcal{D}}f_{\mathcal{S}}(\boldsymbol{x}_{\mathcal{S}}).$
Each term $f_{\mathcal{S}}(\boldsymbol{x}_{\mathcal{S}})$ captures the contribution of a feature subset $\mathcal{S}$ to the function output, reflecting both individual effects and higher-order interactions. This decomposition provides a principled framework for analyzing how input variables influence model predictions. By avoiding the need to estimate expectations and explicitly modeling the influence of feature combinations, it has recently gained significant attention in the explainability literature [23, 31, 32]. Bordt and von Luxburg [23] connect functional decomposition with the Shapley value by defining the value function in terms of the $f_{\mathcal{S}}$ components, thereby extending SHAP to account for both main effects and higher-order interactions. Fumagalli et al. [31] further explore this connection by analyzing various local and global games derived from functional decompositions, offering a unified perspective on local interpretability (feature attributions for individual predictions) and global interpretability (aggregate contributions across the dataset).
In the next section, we build on this line of work by defining the value function via functional decomposition and showing that, for product-kernel methods, this value function arises naturally. We then introduce a recursive formulation that enables the exact computation of Shapley values in polynomial time—significantly improving over the exponential complexity of standard approaches.
3 Exact Shapley value computation for product-kernel learning methods
@@
3.1 Functional decomposition and game-theoretic analysis
Product-kernel methods rely on kernel functions to capture complex relationships between input features and output. A kernel-based decision function is generally expressed as $f(\boldsymbol{x})=\sum_{i=1}^{n}\alpha_{i}k(\boldsymbol{x},\boldsymbol{x}^{(i)})$, where $k$ is a product kernel function, and $\alpha_{i}$ are the learned coefficients associated with the model. The product kernel $k$ can be expressed as products of base kernels $k_{j},j\in\mathcal{D}$:
@@ 3.1 Functional decomposition and game-theoretic analysis
Product-kernel methods are widely used in machine learning for their simplicity and effectiveness in modeling similarities in high-dimensional data—by designing a kernel for each feature and then multiplying them together [33]. They also come with strong theoretical guarantees: if the base kernels are universal—i.e., capable of approximating any continuous function defined on the marginal input—then the product kernel inherits this property. Specifically, Szabó and Sriperumbudur [22] show that product kernels preserve universality and can approximate any continuous function on the whole domain arbitrarily well. Some well-known kernels, such as the radial basis function (RBF) with an isotropic or anisotropic bandwidth, belong to the family of product kernels, i.e., $k(\boldsymbol{x},\boldsymbol{x}^{\prime})=\exp(-\|\boldsymbol{x}-\boldsymbol{x}^{\prime}\|^{2}/2\sigma^{2})=\prod_{j=1}^{d}\exp(-(x_{j}-x^{\prime}_{j})^{2}/2\sigma^{2})$ (see Appendix A for more details on product kernels).
For a product kernel, the decision function naturally induces a functional decomposition.

### v2

@@ Abstract
Kernel methods are widely used in machine learning due to their flexibility and expressiveness. However, their black-box nature poses significant challenges to interpretability, limiting their adoption in high-stakes applications. Shapley value-based feature attribution techniques, such as SHAP and kernel method-specific adaptation like RKHS-SHAP, offer a promising path toward explainability. Yet, computing exact Shapley values is generally intractable, leading existing methods to rely on approximations and thereby incur unavoidable error. In this work, we introduce PKeX-Shapley, a novel algorithm that utilizes the multiplicative structure of product kernels to enable the exact computation of Shapley values in polynomial time. The core of our approach is a new value function, the functional baseline value function, specifically designed for product-kernel models. This value function removes the influence of a feature subset by setting its functional component to the least informative state. Crucially, it allows a recursive thus efficient computation of Shapley values in polynomial time. As an important additional contribution, we show that our framework extends beyond predictive modeling to statistical inference. In particular, it generalizes to popular kernel-based discrepancy measures such as the Maximum Mean Discrepancy (MMD) and the Hilbert–Schmidt Independence Criterion (HSIC), thereby providing new tools for interpretable statistical inference.
Keywords: Shapley value, Kernel Methods, Product Kernel
Introduction
@@
Shapley values (Shapley, 1953), a solution concept originating from cooperative game theory, offer a principled, axiomatic framework for feature attribution in machine learning (ML) (Lundberg and Lee, 2017; Covert et al., 2020). Thanks to their rigorous axiomatic foundation, there has been a widespread adoption within the explainable AI community (Sundararajan and Najmi, 2020). They allow a model’s output—such as predictions or losses—to be fairly distributed across input features based on their individual and joint contributions. Consequently, several algorithms have been proposed to estimate Shapley values under different modelling scenarios, ranging from model-agnostic methods like Kernel SHAP (Lundberg and Lee, 2017) to model-specific methods that leverage structural properties to improve statistical or computational efficiency. Well-known examples of the latter include Tree SHAP for tree-based models (Lundberg et al., 2020), GPSHAP for Gaussian processes (Chau et al., 2024), Deep SHAP for deep networks (Lundberg and Lee, 2017), and, most relevant to our work, RKHS-SHAP for kernel methods (Chau et al., 2022). Kernel methods are particularly notable—not only for their use in prediction tasks but also in a broad range of statistical inference problems, including measuring distributional closeness (Gretton et al., 2006; Naslidnyk et al., 2025), two-sample testing (Schrab, 2025; Chau et al., 2025), goodness-of-fit testing (Chwialkowski et al., 2016; Liu et al., 2016), independence testing (Albert et al., 2022), causal inference and discovery (Mitrovic et al., 2018; Chau et al., 2021; Sejdinovic, 2024), among others. Hence, as kernel methods gain widespread adoption in high-stakes applications for their flexibility and expressive power, the need for interpretability has become increasingly vital.
@@
Despite their game-theoretic foundation, Shapley values face two major challenges when applied to machine learning. The first is defining and estimating a suitable value function that quantifies the contribution of a coalition of features. Ideally, this function should capture the model’s behavior when the complementary features are absent. A natural approach is to retrain the model on each subset and use the resulting prediction as the value (Lipovetsky and Conklin, 2001), but this is computationally infeasible due to the exponential number of retrainings required. A more common alternative simulates feature absence through marginal or conditional expectations of the model’s output, given the fixed subset (see Sundararajan and Najmi (2020) for other formulations). However, estimating these expectations reliably is difficult, as it often involves density estimation or simplifying assumptions such as feature independence—assumptions that can lead to misleading and unfaithful explanations (Kumar et al., 2020). Chau et al. (2022) tackle this problem in the context of kernel methods by exploiting the structure of reproducing kernel Hilbert spaces (RKHS) and using kernel distributional embeddings (Muandet et al., 2017) to estimate value functions nonparametrically, thereby avoiding density estimation and independence assumptions.
@@
Once a value function is fixed, the second challenge lies in efficiently estimating the Shapley values themselves. Exact computation requires evaluating the value function over all $2^{d}$ possible feature subsets for $d$ features, which is computationally demanding. To alleviate this, approximation techniques such as Monte Carlo sampling and regression-based methods are widely used, relying on a smaller set of evaluations. While these approaches reduce computational costs, they inevitably introduce estimation errors that grow with both sample size and feature dimensionality (Kumar et al., 2020). In some cases, model-specific structures can be exploited for efficiency—for example, Tree SHAP leverages tree decompositions for exact polynomial-time computation. By contrast, RKHS-SHAP improves the statistical estimation of the value function but still relies on regression-based approximations for the Shapley values themselves, and thus inherits the same computational limitations.
Figure 1: Execution time of brute-force and our PKeX-Shapley with a 300s budget (see details in Appendix H.2).
To design an efficient Shapley value–based attribution algorithm for kernel methods, we focus on a subclass that employs product kernels—referred to as product-kernel methods—and introduce PKeX-Shapley (Product-Kernel-based eXact Shapley attribution). Product kernels are particularly attractive as they are simple to implement while retaining strong theoretical guarantees. For instance, the product of universal kernels remains universal, so the resulting RKHS can approximate any continuous function provided that each component kernel is sufficiently expressive (Szabó and Sriperumbudur, 2018). Although this focus narrows the scope compared with Chau et al. (2022), which considers generic kernels, the additional structure enables us to directly address the two challenges outlined above. Specifically, we introduce a new value function, called the functional baseline value function, which leverages the multiplicative structure of product kernels. Conceptually, it parallels baseline value functions: rather than setting features to fixed baseline inputs, we set the corresponding kernel components to their least informative state, thereby eliminating their influence. This design admits a natural interpretation as the optimal orthogonal projection of the predictive function onto a function class restricted to a given feature subset. Crucially, this formulation enables a recursive algorithm for computing exact Shapley values in polynomial time (see Figure 1 for a runtime comparison with naïve computation). This yields substantial efficiency gains without relying on approximation techniques such as sampling or regression-based estimation.
As an additional, independent contribution, we demonstrate that our framework extends beyond kernel-based predictive models. In particular, we show how our method applies to kernel-based statistical discrepancies, including the Maximum Mean Discrepancy (MMD) and Hilbert–Schmidt Independence Criterion (HSIC), thereby enabling interpretable statistical inference. A Python implementation of our method is publicly available (pke, 2025).
Background on Shapley values
@@
Notation. Let $\mathcal{D}$ denote the set of $d$ features, and $2^{\mathcal{D}}$ its power set. The training set $\{(\boldsymbol{x}^{(i)},y^{(i)})\}_{i=1}^{n}$ consists of $n$ samples, where $\boldsymbol{x}^{(i)}\in\mathbb{R}^{d}$ and $y^{(i)}\in\mathbb{R}$ (or a discrete label set for classification tasks). Let $\mathbf{X}\in\mathbb{R}^{n\times d}$ be the matrix of features, and $\mathbf{X}_{\mathcal{S}}\in\mathbb{R}^{n\times s}$ the submatrix restricted to features in subset $\mathcal{S}\subseteq\mathcal{D}$, and we write $\mathbf{X}_{j}:=\mathbf{X}_{\{j\}}$. We use capital letters for random variables, bold capital letters for matrices, calligraphic letters for sets, and bold lowercase letters for vectors. The restriction of a vector $\boldsymbol{x}$ to features in $\mathcal{S}$ is denoted by $\boldsymbol{x}_{\mathcal{S}}$. The elementwise product is denoted by $\odot$, and expectation by $\mathbb{E}$. A symmetric positive (semi-)definite kernel function over $\mathcal{D}$ is denoted by $k$, and its restriction to subset $\mathcal{S}$ by $k_{\mathcal{S}}$. The corresponding kernel matrices are denoted by $\mathbf{K}$ and $\mathbf{K}_{\mathcal{S}}$, respectively. All proofs are provided in Appendix D.
@@
Shapley value. The Shapley value (Shapley, 1953) is a widely used game-theoretic method for feature attribution in predictive models. It assigns importance to each input feature by averaging its marginal contribution across all possible feature subsets, with weights determined by coalition size. This construction uniquely satisfies several fairness axioms, making it a principled approach to distributing a model’s output among its features. Specifically, given a value function $v:2^{\mathcal{D}}\to\mathbb{R}$, which quantifies the contribution of a feature subset, the Shapley value for a feature $j$ is defined as
@@
where $\mu(s)=\nicefrac{{s!(d-s-1)!}}{{d!}}$ and $v(\mathcal{S}\cup\{j\})-v(\mathcal{S})$ measures the marginal contribution of feature $j$ when added to subset $\mathcal{S}$. The Shapley value is the unique solution satisfying four axioms of efficiency, null player, symmetry, and linearity for any given value function (Shapley, 1953). However, in the context of explainability, defining this value function vv—which determines how subsets of features contribute to the model’s output—is a non-trivial design choice. Several formulations have been proposed (Sundararajan and Najmi, 2020), but two are widely adopted: the interventional value function $v_{\boldsymbol{x}}(\mathcal{S})=\mathbb{E}_{X_{\mathcal{D}\setminus\mathcal{S}}|X_{\mathcal{S}}=\boldsymbol{x}_{\mathcal{S}}}[f(\boldsymbol{x}_{\mathcal{S}},X_{\mathcal{D}\setminus\mathcal{S}})]$, which replaces missing features with samples from their marginal distributions (Janzing et al., 2020), and the observational value function $v_{\boldsymbol{x}}(\mathcal{S})=\mathbb{E}_{X_{\mathcal{D}\setminus\mathcal{S}}}[f(X)\,|\,X_{\mathcal{S}}=\boldsymbol{x}_{\mathcal{S}}]$, which imputes them using the conditional distribution given observed features (Lundberg and Lee, 2017). While the observational approach preserves feature dependencies, it requires estimating complex conditional expectations, making it computationally demanding. In practice, both approaches are often implemented by sampling from marginal or conditional distributions, requiring repeated model evaluations that become costly for large models or high-dimensional inputs (Sundararajan and Najmi, 2020). Moreover, these estimates typically rely on a finite set of background samples (commonly drawn from the training set), which can substantially influence the resulting explanations (Molnar, 2023, Chapter 21).
Exact Shapley value for product-kernel learning methods
@@
3.1 A new functional baseline value function for product-kernel methods
Product-kernel methods rely on kernel functions to capture complex relationships between input features and output. A kernel-based decision function is generally expressed as
$f(\boldsymbol{x})=\sum_{i=1}^{n}\alpha_{i}k(\boldsymbol{x},\boldsymbol{x}^{(i)})=\boldsymbol{\alpha}^{\top}k(\mathbf{X},\boldsymbol{x}){}$
where $k$ is a product kernel function, and $\alpha_{i}$ are the learned coefficients associated with the model. Recall that the product kernel $k$ can be expressed as products of base kernels $k_{j},j\in\mathcal{D}$:
@@ 3.1 A new functional baseline value function for product-kernel methods
Product-kernel methods are widely used in machine learning for their simplicity and effectiveness in modeling similarities in high-dimensional data—by designing a kernel for each feature and then multiplying them together (Gardner et al., 2018). They also come with strong theoretical guarantees: if the base kernels are universal—i.e., capable of approximating any continuous function defined on the marginal input—then the product kernel inherits this property (Szabó and Sriperumbudur, 2018). Some well-known kernels, such as the radial basis function (RBF) with an isotropic or anisotropic bandwidth, belong to the family of product kernels, i.e., $k(\boldsymbol{x},\boldsymbol{x}^{\prime})=\exp(-\|\boldsymbol{x}-\boldsymbol{x}^{\prime}\|^{2}/2\sigma^{2})=\prod_{j=1}^{d}\exp(-(x_{j}-x^{\prime}_{j})^{2}/2\sigma^{2})$ (see Appendix A for more details on product kernels).
We now introduce our functional baseline value function, motivated by the multiplicative structure of product kernels. In standard settings, interventional value functions remove the effect of a feature by replacing it with its average contribution, thereby isolating its marginal impact. More generally, this can be viewed as a baseline value function (Sundararajan and Najmi, 2020, Sec. 2.3), where features are set to a non-informative state, such as zero vectors in image data. In our setting, however, the functional form of the predictive model is directly accessible. We therefore remove a feature’s influence by factoring out its corresponding functional component and setting it to one—the least informative contribution in the multiplicative structure. This design principle yields the following definition for product-kernel methods.
Definition 1.
Let $\boldsymbol{\alpha}$ be the coefficients of the function $f$ constructed using the product-kernel method. The functional baseline value function for an input $\boldsymbol{x}$, a feature coalition $\mathcal{S}$, and coefficients $\boldsymbol{\alpha}$ is
$\nu_{\boldsymbol{x}}(\mathcal{S})=\boldsymbol{\alpha}^{\top}k_{\mathcal{S}}(\mathbf{X}_{\mathcal{S}},\boldsymbol{x}_{\mathcal{S}}).$ (3)
Notably, computing this value function requires only linear time, making it highly efficient. A detailed comparison with baseline and interventional value functions is provided in Appendix B. Interestingly, our functional baseline value function can also be understood as the orthogonal projection of $f$ onto an RKHS subspace of functions depending only on features in $\mathcal{S}$. We state the formal proposition below and defer the full discussion along with the proof in Appendix B.3.
Proposition 2.
$\nu_{\boldsymbol{x}}(\mathcal{S})$ is the orthogonal projection of $f$ to $\mathbb{H}_{\mathcal{S}}:=\{g\in\mathcal{H}:g(\boldsymbol{x})=g(\boldsymbol{x}^{\prime})\text{ when }\boldsymbol{x}_{\mathcal{S}}=\boldsymbol{x}^{\prime}_{\mathcal{S}}\}$, the subspace of functions in $\mathcal{H}$ depending only on features in $\mathcal{S}$.
We emphasize that the choice of a value function determines how feature subset importance is measured, and is therefore inherently subjective. Our introduction of the functional baseline value function is motivated not only by its conceptual alignment with existing formulations, but also by the substantial computational benefits it enables, as we will demonstrate in the following subsection. We do not claim that this value function is intrinsically superior in capturing feature importance; however, we demonstrate in Experiment #2 (Section 5.1) that its empirical performance in recovering influential features is promising and on par with other established baselines.
3.2 A recursive formulation for computing Shapley values
Having motivated our functional baseline value function, we now demonstrate the computational gain we obtain from this formulation.